The sum of two numbers is 52, and their difference is 6. What are the numbers?
Answer by mananth(12270)

let one number be x and the other be y

x+y=52
x-y=6
add the two equations to eliminate y
2x=58
x= 58/2
x=29
y= 52-29= 23